OpenStudy (anonymous):

Simplify f(x)+g(x) when f(x)=-6/(x+3) and g(x)= 2/x

OpenStudy (anonymous):

This is my last one! Would you mind walking me through this one last time? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

In order to add the fractions, you need to make all the denominators equal to the LCD. Once you do that, you can add the fractions. \[\Large f(x)+g(x)\] \[\Large -\frac{6}{x+3}+\frac{2}{x}\] \[\Large -\frac{6x}{x(x+3)}+\frac{2}{x}\] \[\Large -\frac{6x}{x(x+3)}+\frac{2(x+3)}{x(x+3)}\] \[\Large -\frac{6x}{x(x+3)}+\frac{2x+6}{x(x+3)}\] \[\Large \frac{-6x+2x+6}{x(x+3)}\] \[\Large \frac{-4x+6}{x(x+3)}\] \[\Large \frac{-4x+6}{x^2+3x}\] So \[\Large f(x)+g(x)=\frac{-4x+6}{x^2+3x}\]

OpenStudy (anonymous):

Thank you so incredibly much again! You have no idea how much I appreciate you walking me through this stuff, thank you =)

jimthompson5910 (jim_thompson5910):

you're very welcome :)